# Crossing Numbers of Beyond Planar Graphs Re-revisited: A Framework Approach

### v1

@@ 1 Introduction
We are interested in the minimum number of crossings within any beyond-planarity restricted drawing. Intuitively speaking, a beyond planarity concept is a way to (try to) formalize our understanding of what aspects constitute a readable drawing. Still, within these restrictions it is natural to ask for a drawing with the least number of crossings. Alternatively, we can ask how much it “costs” in terms of crossings to follow a certain drawing paradigm.
Formally, let $\mathfrak{B}(k)$ denote any beyond planarity concept, where we omit the parameter $k$ for parameterless beyond planarities. Then, $\mathcal{G}_{\mathfrak{B}(k)}$ ($\mathcal{G}_{\mathfrak{B}(k)}(n)$) denotes the set of graphs (on $n$ vertices, respectively) allowing a $\mathfrak{B}(k)$-drawing. The $\mathfrak{B}(k)$-crossing number $\mathrm{cr}_{\mathfrak{B}(k)}(G)$ for a graph $G\in\mathcal{G}_{\mathfrak{B}(k)}$ is the least number of crossings over all $\mathfrak{B}(k)$-drawings of $G$. The crossing ratio $\varrho_{\mathfrak{B}(k)}(n)$ of $\mathfrak{B}(k)$ is the largest attainable ratio between the $\mathfrak{B}(k)$-crossing number and the (normal) crossing number, over all $n$-vertex $\mathfrak{B}(k)$-graphs:
@@
Crossing ratios of beyond planar concepts have for the first time been explicitly considered in [CKMV19], where linear lower bounds $\Omega(n)$ where established for 1-planar, $k$-quasi-planar, and (weakly) fan-planar graphs. In all but the first case, the upper bounds where (at least) quadratic in $n$, and it was conjectured in [CKMV19] that the real lower bounds should rather be $\Omega(n^{2})$. Later, [BPS21] generalized the results to $k$-planarity and seven more beyond planarity concepts. However, none of the provided bounds were tight except for planarly-connected and straight-line RAC. In both publications, each planarity measure needs a very specific construction and intricate specialized arguments to prove the lower bounds (which tend to always be the most complicated part in crossing number proofs). Furthermore, both only consider simple drawings, when establishing upper bounds on the crossing ratio. Simple and non-simple crossing ratios may differ, if there are $\mathfrak{B}(k)$-graphs that only allow non-simple $\mathfrak{B}(k)$-drawings. Our bounds below work with and without the simplicity assumption.
@@
In this paper, we propose a general framework that, when applicable, simplifies crossing ratio proofs down to only a couple of $\mathfrak{B}(k)$-specific lines (10–20 lines for the lower bound). In contrast to previous schemes, this framework is able to prove tight bounds for all considered beyond planarities, and our proofs further show that the crossing ratio is already achieved between any two subsequent parameterizations $k$ and $k+1$, in all considered parametrized concepts. Table 1 summarizes our results. A key idea lies in the simplification of the necessary lower bound arguments, by turning them into counting arguments over a set of (sufficiently) disjoint Kuratowski-subdivisions (see Section 2). In particular, this allows us to avoid any intricate discussions about alternative drawing possibilities, the simplicity of drawings, etc. We present the proof framework in Section 3, and showcase its versatility and strength for various beyond planarity concepts in LABEL:sec:use.
Table 1: Overview on our results on the crossing ratios of beyond planarities for $n$-vertex graphs. Formally, [CKMV19] considers weakly fan-planar, but their proof works for all 4 variants. The column “$\overline{\varrho}$?” denotes whether we also show the (same) rectilinear crossing ratios, see LABEL:sec:rectilinear. The -marked previous upper bounds need no explicit discussion of non-simple drawings by the nature of the beyond planarity concept. All our bounds work with and without the simplicity assumption.
beyond planarity previous best our results Section $\overline{\varrho}$?
@@
$k$-apex-planar $\Omega(n/k)\cap\mathcal{O}(k^{2}n^{2}+k^{4})$ [BPS21] $\Theta(n^{2}/k)$ LABEL:sec:kapex
skewness-$k$ $\Omega(n/k)\cap\mathcal{O}(kn+k^{2})$ [BPS21] $\Theta(n)$ LABEL:sec:kapex
@@ 2 Preliminaries
$\varrho_{\mathfrak{B}(k)}(n)\leq\max\left\{\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)\ ,\ \sup_{G\in\mathcal{G}^{\mathrm{dense}}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}\right\},$
@@
    If $\mathrm{cr}_{\mathfrak{B}(k)}(G)\in\mathcal{O}(km)$, then $\varrho_{\mathfrak{B}(k)}(n)\in{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+km\cdot\frac{n^{2}}{m^{3}})\subseteq}\;\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+k).$
@@
The crossing ratio $\varrho_{\mathfrak{B}(k)}(n)$ relates the $\mathfrak{B}(k)$-restricted crossing number to the normal one. It is natural to ask, how small a parameter $\Delta\geq 1$ can be chosen such that $\varrho_{\mathfrak{B}(k)}^{\Delta}(n)\coloneqq\sup_{G\in\mathcal{G}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}_{\mathfrak{B}(k+\Delta)}(G)}$ attains the same asymptotic bound. We say $\varrho_{\mathfrak{B}(k)}(n)$ is $\Delta$-jumping if $\varrho_{\mathfrak{B}(k)}^{\Delta}(n)\in\Omega(\varrho_{\mathfrak{B}(k)}(n))$. In LABEL:sec:use, we will always observe 1-jumping as our upper bounds on the normal crossing numbers are attained by $\mathfrak{B}(k+1)$-drawings.

### v2

@@ 1 Introduction
We are interested in the minimum number of crossings within any beyond planarity restricted drawing. Intuitively speaking, a beyond planarity concept is a way to (try to) formalize our understanding of what aspects constitute a readable drawing. Still, within these restrictions it is natural to ask for a drawing with the least number of crossings. Alternatively, we can ask how much it “costs” in terms of crossings to follow a certain drawing paradigm.
Formally, let $\mathfrak{B}(k)$ denote any beyond planarity concept, where we omit the parameter $k$ for parameterless beyond planarity concepts. Then, $\mathcal{G}_{\mathfrak{B}(k)}$ ($\mathcal{G}_{\mathfrak{B}(k)}(n)$) denotes the set of graphs (on $n$ vertices, respectively) allowing a $\mathfrak{B}(k)$-drawing. The $\mathfrak{B}(k)$-crossing number $\mathrm{cr}_{\mathfrak{B}(k)}(G)$ for a graph $G\in\mathcal{G}_{\mathfrak{B}(k)}$ is the least number of crossings over all $\mathfrak{B}(k)$-drawings of $G$. The crossing ratio $\varrho_{\mathfrak{B}(k)}(n)$ of $\mathfrak{B}(k)$ is the largest attainable ratio between the $\mathfrak{B}(k)$-crossing number and the (normal) crossing number, over all $n$-vertex $\mathfrak{B}(k)$-graphs:
@@
Crossing ratios of beyond planarity concepts have for the first time been explicitly considered in [CKMV19a, CKMV19], where linear lower bounds $\Omega(n)$ where established for 1-planar, $k$-quasi-planar, and (weakly) fan-planar graphs. In all but the first case, the upper bounds where (at least) quadratic in $n$, and it was conjectured in [CKMV19] that the real lower bounds should rather be $\Omega(n^{2})$. Later, [BPS21] generalized the results to $k$-planarity and seven more beyond planarity concepts. However, none of the provided bounds were tight except for planarly-connected and straight-line RAC. In both publications, each planarity measure needs a very specific construction and intricate specialized arguments to prove the lower bounds (which tend to always be the most complicated part in crossing number proofs). Furthermore, both only consider simple drawings, when establishing upper bounds on the crossing ratio. Simple and non-simple crossing ratios may differ, if there are $\mathfrak{B}(k)$-graphs that only allow non-simple $\mathfrak{B}(k)$-drawings. Our bounds below work with and without the simplicity assumption.
@@
In this paper, we propose a general framework that, when applicable, simplifies crossing ratio proofs down to only a couple of $\mathfrak{B}(k)$-specific lines (10–20 lines for the lower bound). In contrast to previous schemes, this framework is able to prove asymptotically tight bounds for all considered beyond planarity concepts, and our proofs further show that the crossing ratio is already achieved between any two subsequent parameterizations $k$ and $k+1$, in all considered parametrized concepts. Table 1 summarizes our results. A key idea lies in the simplification of the necessary lower bound arguments, by turning them into counting arguments over a set of (sufficiently) disjoint Kuratowski-subdivisions (see Section 2). In particular, this allows us to avoid any intricate discussions about alternative drawing possibilities, the simplicity of drawings, etc. We present the proof framework in Section 3, and showcase its versatility and strength for various beyond planarity concepts in LABEL:sec:use.
Table 1: Overview on our results on the crossing ratios of beyond planarity concepts for $n$-vertex graphs. Formally, [CKMV19] considers weakly fan-planar, but their proof works for all 4 variants. The column “$\overline{\varrho}$?” denotes whether we also show the (same) rectilinear crossing ratios in LABEL:sec:rectilinear. The -marked previous upper bounds need no explicit discussion of non-simple drawings by the nature of the beyond planarity concept. All our bounds, except the upper bounds for adjacency-crossing and fan-crossing, work with and without the simplicity assumption.
beyond planarity concept previous best our results Section $\overline{\varrho}$?
@@
$k$-apex-planar $\Omega(n/k)\cap\mathcal{O}(k^{2}n^{2})$ [BPS21] $\Theta(n^{2}/k)$ LABEL:sec:kapex
skewness-$k$ $\Omega(n/k)\cap\mathcal{O}(kn)$ [BPS21] $\Theta(n)$ LABEL:sec:kapex
@@ 2 Preliminaries
$\varrho_{\mathfrak{B}(k)}(n)=\max\left\{\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)\ ,\ \sup_{G\in\mathcal{G}^{\mathrm{dense}}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}(G)}\right\},$
@@
    If $\mathrm{cr}_{\mathfrak{B}(k)}(G)\in\mathcal{O}(mk)$, then $\varrho_{\mathfrak{B}(k)}(n)\in{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+mk\cdot\frac{n^{2}}{m^{3}})\subseteq}\;\mathcal{O}(\varrho_{\mathfrak{B}(k)}^{\mathrm{sparse}}(n)+k).$
@@
The crossing ratio $\varrho_{\mathfrak{B}(k)}(n)$ relates the $\mathfrak{B}(k)$-restricted crossing number to the normal one. It is natural to ask, how small a parameter $\Delta\geq 1$ can be chosen such that $\varrho_{\mathfrak{B}(k)}^{\Delta}(n)\coloneqq\sup_{G\in\mathcal{G}_{\mathfrak{B}(k)}(n)}\frac{\mathrm{cr}_{\mathfrak{B}(k)}(G)}{\mathrm{cr}_{\mathfrak{B}(k+\Delta)}(G)}$ attains the same asymptotic bound. We say $\varrho_{\mathfrak{B}(k)}(n)$ is $\Delta$-jumping if $\varrho_{\mathfrak{B}(k)}^{\Delta}(n)\in\Omega(\varrho_{\mathfrak{B}(k)}(n))$. For all concepts in LABEL:sec:use except $k$-gap-planarity (LABEL:sec:kgappl), we will observe 1-jumping as our upper bounds on the normal crossing numbers are already attained by $\mathfrak{B}(k+1)$-drawings.